8. A vertical composite fluid column whose upper end is open to the atmosphere is composed of 20

inches of mercury (S.G = 13.45), 30 inches of water (density = 62.4 lb m /ft 3 ) and 35 inches of oil (S.G =

0.825). Determine the absolute pressure in psia: a) at the base of the column, b) at the oil-water

interface, c) at the water-mercury interface.


1
Expert's answer
2021-09-10T00:03:08-0400

Part a

Let P3abs = absolute pressure at the bottom of column "= P_3 + Patm"

"P_3abs = P_3 + Patm = 81698 Pa + 101325 Pa = 183023 Pa\\

P_3abs = 183023 Pa = (183023 * 0.000145038) psai = 26.545 psai\\

P_3abs = 26.5 psai"


Part b

Let "P_1" = gauge pressure at oil-water interface = "d_3 g h_3"

"P_1 = d_3 g h_3 = 825 x 9.81x 0.889 = 7195 Pa\\

P_1 = 7195 Pa"

Let "P_2" = gauge pressure at water-mercury interface ="P1 + d_2 g h_2"

"P_2 = P1 + d_2 g h_2 = 7195 Pa + (1000 x 9.81 x 0.762 ) = 14670 Pa\\

P_2 = 14670 Pa"

Let P3 = gauge pressure at the bottom of column = "P_2 + d_1 g h_1"

"P_3 = P_2 + d_1 g h_1 =14670 Pa + (13450 x 9.81 x 0.508 ) = 81698 Pa\\

P_3 = 81698 Pa"

Let P1abs = absolute pressure at oil-water interface = P1 + Patm

"P_{1}abs = P_1 + P_{atm} = 7195 Pa + 101325 Pa = 108520 Pa\\

P_{1}abs = 108520 Pa = (108520 * 0.000145038) psai = 15.739 psai\\

P_1abs = 15.7 psai"


Part c

Let P2abs = absolute pressure at water-mercury interface = "P_2 + Patm"

"P_2abs = P_2 + Patm = 14670 Pa + 101325 Pa = 115995 Pa\\

P_2abs = 115995 Pa = (115995 * 0.000145038) psai = 16.824 psai\\

P_2abs = 16.8 psai"